A sample of lawn fertilizer was analyzed by back titration to determine its ammonium ion content. The sample was boiled in NaOH solution. The reaction occurred according to the equation

NH4+(aq) + NaOH(aq)


1
Expert's answer
2020-07-30T07:09:11-0400

General Description and Properties.

Ammonium ion is formed by the reaction between acids and aqueous ammonia:


NH3 (aq)+ H+ (aq) = NH > 4 + (aq)


The ammonium ion behaves chemically like the ions of the alkali metals, particularly potassium ion, which is almost the same size. All ammonium salts are white and soluble.

Characteristic reactions of NH₄⁺Sodium Hydroxide

Addition of concentrated hydroxide ion solutions evolves NH3 gas:


NH4+ (aq) + OH- (aq) = NH3 (aq) + H2O (l)


The ammonia gas turns moistened red litmus paper blue. It can also be detected by its characteristic odor. This reaction serves as a confirmatory test for NH4+